OpenStudy (anonymous):

What is the solution of the equation? 6x – 3 = 5x – 5 A. –4 B. –2 C. 0 D. –1

OpenStudy (anonymous):

6x – 3 = 5x – 5 we want to solve for x that means that we want all the x's on one side. Where do you want it on the left side or right side?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

right side

OpenStudy (anonymous):

great we need to subtract 6x to both sides then 6x – 3 = 5x – 5 6x-6x – 3 = 5x-6x – 5 -3 =-x -5 Now we need to add 5 to both sides in order to get rid of that -5 on the right side. What is the equation then?

OpenStudy (anonymous):

to be honest the letters are confusing me i suck in math this is for my son

OpenStudy (anonymous):

Can you let your son try.

OpenStudy (anonymous):

yes i can i just dont want him to know how bad i am in math so i was trying to help him on the slick side this is a real gut buster here for me to not be able to help him

OpenStudy (anonymous):

6x - 3 = 5x - 5 First you can subtract 5x from both sides of the equation to get all the 'x' terms on the same side: 6x - 3 = 5x - 5 6x - 3 - 5x = 5x - 5x - 5 x - 3 = - 5 Now we can add 3 to both sides of the equation to find out what 'x' equals: x - 3 = - 5 x - 3 + 3 = - 5 + 3 x = - 2 , this is the solution to the equation.